OpenStudy (anonymous):

Divide... 4x^3-8x^2+19 / x-2 please

OpenStudy (anonymous):

you can use the long method or the synthetic division...

OpenStudy (anonymous):

let's use synthetic division: REMEMBER to solve for x of the divisor (in this case the divisor is x-2): wherein x = 2 (x-2=0) (x=2) 

OpenStudy (anonymous):

give your conclusion...@KE8717504

OpenStudy (anonymous):

no idea i hate doing these honestly

OpenStudy (anonymous):

hello again

OpenStudy (anonymous):